OpenStudy (anonymous):

log3 2x- 5 log3 y A. log3 (2x- y^5) B. log3 (2x-5y) C. log3 2x/ 5y D. log3 2x/y^5

OpenStudy (anonymous):

lol i forgot how to do this

OpenStudy (anonymous):

thats okay. me 2!

OpenStudy (danjs):

\[a * \log _{b}x = \log _{b} x ^{a}\]

OpenStudy (danjs):

and one more thing

Nnesha (nnesha):

so subtraction become what ?? opposite of last one hope you remember

OpenStudy (danjs):

\[\log(x) - \log(y) = \log \frac{ x }{ y }\]

OpenStudy (anonymous):

hope these guys can help ya

OpenStudy (danjs):

Those are the 2 rules you need to use

OpenStudy (anonymous):

+

Nnesha (nnesha):

nope subtraction -----> division addition ------> multiplication

OpenStudy (danjs):

\[\log _{3}(2x) - \log _{3}(y^5)\]

OpenStudy (anonymous):

oh

OpenStudy (danjs):

that is the first step, only one left to do

Nnesha (nnesha):

this one is opposite to last one last time was expand log now this time condense

Nnesha (nnesha):

 power become multiplication of log so first change that 5 log3 y